OpenStudy (anonymous):

HELP ASAP WILL MEDAL HELP ASAP WILL MEDAL Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (zpupster):

slope formula m=y2-y1/x2-x1 m = slope y2=0 y1=4 x2=2 x1=-6 plugging in 0-4/2-(-6)= slope point-slope form y-y1=m(x-x1) now you can pick either (x,y) point because both are on the line but it looks like your answer are (-6,4) pair y-4=m(x-(-6)

OpenStudy (anonymous):

is it a or b?

OpenStudy (zpupster):

well what is your slope??

OpenStudy (anonymous):

so b

OpenStudy (zpupster):

look at this part 0-4/2-(-6)= slope just do the math or do you need help?

OpenStudy (anonymous):

i got it Thanks!

OpenStudy (zpupster):

ok